# Do Large Language Models Grasp The Grammar? Evidence from Grammar-Book-Guided Probing in Luxembourgish

### v1

@@ Abstract
Grammar refers to the system of rules that governs the structural organization and the semantic relations among linguistic units such as sentences, phrases, and words within a given language. In natural language processing, there remains a notable scarcity of grammar-focused evaluation protocols, a gap that is even more pronounced for low-resource languages. Moreover, the extent to which large language models genuinely comprehend grammatical structure especially the mapping between syntactic structures and meanings remains under debate. To investigate this issue, we propose a Grammar-Book–Guided evaluation pipeline intended to provide a systematic and generalizable framework for grammar evaluation consisting of four key stages, and in this work we take Luxembourgish as a case study. The results show a weak positive correlation between translation performance and grammatical understanding, indicating that strong translations do not necessarily imply deep grammatical competence. Larger models perform well overall due to their semantic strength but remain weak in morphology and syntax, struggling particularly with Minimal Pair tasks, while strong reasoning ability offers a promising way to enhance their grammatical understanding. The code and prompts are open-sourced at : https://anonymous.4open.science/r/GBGP-504E.
@@ 1. Introduction
In recent years, Large Language Models (LLMs) have transformed the field of Natural Language Processing (NLP), achieving remarkable progress across a variety of tasks such as text generation and translation Minaee et al. (2024). Their ability to produce coherent, and contextually relevant text has raised important discussions about whether such models merely mimic language patterns or demonstrate a deeper understanding of linguistic structures. Furthermore, while much of the early focus has been on their fluency and performance in high-resource languages such as English, French, or German, less attention has been paid to their grammatical competence in low-resource languages Li et al. (2024).
Grammar Challenges Machine translation, in particular, is an area where LLMs are widely applied and evaluated. Success is often measured in terms of fluency and semantic adequacy compared to the reference (golden) translation; however, these metrics do not necessarily capture whether the model adheres to the explicit grammatical rules of the target languageCaglayan et al. (2020). For widely used languages with abundant training data, grammar is often implicitly learned through statistical exposure. However, the challenge becomes much more pronounced for low-resource languages, which have limited digital representation and many corner cases that are especially difficult to capture during trainingLi et al. ; Zhong et al. (2024); Cahyawijaya et al. (2024). In such cases, an LLM’s ability to generate correct translations may not imply genuine grammatical understanding. In this sense, Luxembourgish presents a compelling case study in this context. As a low-resource language with fewer available corpora compared to its neighboring languages (German and French), Luxembourgish can pose difficult challenges on both translation and language understandingLothritz et al. (2022).
@@ 2. Related Work
BERT Moment in Translation LLM remain controversial regarding whether they truly understand semantics or merely mimic human speech; a potential attitude that makes one wonder whether an LLM does not amount to being just a “stochastic parrot” Bender et al. (2021). Despite these debates, their performance in language translation has become unparalleled, vastly exceeding many previous non-transformer and even encoder–decoder models Tiedemann et al. ; Team et al. (2022a). This development constitutes a genuine “BERT moment” for translation research. Previous non-conversational models showed limited logical reasoning, and their multilingual ability was mostly attributed to abstract mechanisms like semantic mapping. In contrast, LLMs mark a major change: by asking targeted questions and running controlled comparisons, one can directly test how they understand and apply grammatical rules, giving clearer, verifiable evidence of their cross lingual abilities.
@@ 3.1. Mindsets - Why Grammar Book?
Translation enhanced by Grammar Our work is inspired by "Machine Translation from One Book" which demonstrates that it is possible to translate a language with virtually no web data (Kalamang) using a grammar book as the primary resource Tanzer et al. (2023). However, Aycock et al. (2024) shows that the translation performance in this setting is primarily attributable to the parallel example sentences embedded within the grammar book, rather than the grammatical descriptions themselves. This raises the question: instead of using the grammar book for translation, can we leverage a complete grammar book to simply evaluate the grammatical understanding capabilities of LLMs? Furthermore, open-source, grammar-focused datasets remain scarce and often exhibit incomplete coverage of grammatical phenomena, with many evaluations restricted to a single minimal-pair testing paradigm whose richness is limited.
@@ 3.2. Pipeline Structure
Material Inspector: component which identifies and typologies grammatical phenomena and error patterns from the grammar document;
@@ 3.3. Models and Datasets
Dataset Avg. Char Len ( ± ) # N Topics Ref. G2040 171.63 ± 21.42 2040 Grammars 2023 LONG300 429.47 ± 49.18 300 News 2025 FLORES200 125.61 ± 41.18 1012 Diverse 2022b
@@ 4.2. Statistical Analysis
We conducted Spearman and Pearson correlation analyses on the performance of four the grammatical comprehension test tasks as well as on translation performance, and the results are illustrated in Figure 2. It can be observed that Tasks 1, 2, and 3 exhibit very strong correlations, primarily reflecting the ability to identify grammatical structures and their corresponding sentences. However, these three tasks show relatively low correlations with Task 4, which involves identifying sentences with correct grammatical usage. As a result, there is a weak positive correlation between performance on grammatical comprehension tasks and translation performance; however, it is insufficient to conclude that strong translation implies grammatical understanding.
In addition, we included the relatively smaller Unbabel/wmt22-cometkiwi-da (Big size) model as a supplementary evaluation to the Unbabel/wmt23-cometkiwi-da-xl (Small size) model. We found that the larger model shows higher correlations with spBLEU and ChrF++, while the smaller model’s evaluation scores are less stable. This finding led us to adopt Unbabel/wmt22-cometkiwi-da as our primary evaluation metric.
@@ 4.3. Translation Vs Grammar Understanding?
Partial mastery of grammar We take the Llama family as a case study and observe that scaling from 1B - 3B to 70B parameters improves translation performance from spBLEU to , which yields partial gains on Task 2 and Task 3 (approximately 30%), yet leaves Task 1 essentially unchanged and Task 4 still near chance-level performance. Comparable patterns emerge for the Qwen and Gemma families. Only GPT-5 achieves consistently superior results across all tasks and surpasses chance performance on Task 4. Moreover, interestingly, reasoning model “Qwen3-Next-80B-A3B-Thinking” demonstrate higher performance on most grammar understanding tasks, even outperforming GPT-5, highlighting the crucial role of logical reasoning in grammatical comprehension. Moreover, increasing model size within the same family confers modest benefits to grammatical understanding - likely via enhanced reasoning - yet these gains are markedly smaller than those observed for translation.
@@ Acknowledgements
- Artifex Software, Inc. et al. (2025) Artifex Software, Inc., Jorj X. McKie, and PyMuPDF Contributors. 2025. Pymupdf: Python library for pdf and document processing. https://github.com/pymupdf/PyMuPDF. Version 1.26.5, AGPL-3.0 License, accessed 2025-10-14.
@@
- Aycock et al. (2024) Seth Aycock, David Stap, Di Wu, Christof Monz, and Khalil Sima’an. 2024. Can llms really learn to translate a low-resource language from one grammar book? ArXiv, abs/2409.19151.
- Bender et al. (2021) Emily M. Bender, Timnit Gebru, Angelina McMillan-Major, and Shmargaret Shmitchell. 2021. On the dangers of stochastic parrots: Can language models be too big? In Proceedings of the 2021 ACM Conference on Fairness, Accountability, and Transparency, FAccT ’21, page 610–623, New York, NY, USA. Association for Computing Machinery.
@@
- Caglayan et al. (2020) Ozan Caglayan, Pranava Madhyastha, and Lucia Specia. 2020. Curious case of language generation evaluation metrics: A cautionary tale. arXiv preprint arXiv:2010.13588.
@@
- Cahyawijaya et al. (2024) Samuel Cahyawijaya, Holy Lovenia, and Pascale Fung. 2024. Llms are few-shot in-context low-resource language learners. arXiv preprint arXiv:2403.16512.
- Cui et al. (2025) Cheng Cui, Ting Sun, Manhui Lin, Tingquan Gao, Yubo Zhang, Jiaxuan Liu, Xueqing Wang, Zelun Zhang, Changda Zhou, Hongen Liu, Yue Zhang, Wenyu Lv, Kui Huang, Yichao Zhang, Jing Zhang, Jun Zhang, Yi Liu, Dianhai Yu, and Yanjun Ma. 2025. Paddleocr 3.0 technical report.
- Duan et al. (2025) Xufeng Duan, Zhaoqian Yao, Yunhao Zhang, Shaonan Wang, and Zhenguang G. Cai. 2025. How syntax specialization emerges in language models.
@@
- Gilles (2006) Peter Gilles. 2006. Phonologie der n-Tilgung im Moselfränkischen (’Eifeler Regel’). Ein Beitrag zur dialektologischen Prosodieforschung, pages 29–68.
- Gilles (2023) Peter Gilles. 2023. Luxembourgish.
@@
- Ide et al. (2025) Yusuke Ide, Yuto Nishida, Justin Vasselli, Miyu Oba, Yusuke Sakai, Hidetaka Kamigaito, and Taro Watanabe. 2025. How to make the most of LLMs’ grammatical knowledge for acceptability judgments. In Proceedings of the 2025 Conference of the Nations of the Americas Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), pages 7416–7432, Albuquerque, New Mexico. Association for Computational Linguistics.
@@
- Kudo and Richardson (2018) Taku Kudo and John Richardson. 2018. SentencePiece: A simple and language independent subword tokenizer and detokenizer for neural text processing. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, pages 66–71, Brussels, Belgium. Association for Computational Linguistics.
@@
- Li et al. (2024) Zihao Li, Yucheng Shi, Zirui Liu, Fan Yang, Ninghao Liu, and Mengnan Du. 2024. Quantifying multilingual performance of large language models across languages. CoRR.
@@
- (13) Zihao Li, Yucheng Shi, Zirui Liu, Fan Yang, Ali Payani, Ninghao Liu, and Mengnan Du. Language ranker: A metric for quantifying llm performance across high and low-resource languages. arxiv 2024. arXiv preprint arXiv:2404.11553.
@@
- Lin (2024) Sha Lin. 2024. Evaluating llms’ grammatical error correction performance in learner chinese. PLOS ONE, 19(10):1–18.
- Lothritz et al. (2022) Cedric Lothritz, Bertrand Lebichot, Kevin Allix, Lisa Veiber, Tegawendé François D Assise Bissyande, Jacques Klein, Andrey Boytsov, Anne Goujon, and Clément Lefebvre. 2022. Luxembert: Simple and practical data augmentation in language model pre-training for luxembourgish. In 13th Language Resources and Evaluation Conference (LREC 2022).
- Minaee et al. (2024) Shervin Minaee, Tomas Mikolov, Narjes Nikzad, Meysam Chenaghlu, Richard Socher, Xavier Amatriain, and Jianfeng Gao. 2024. Large language models: A survey. arXiv preprint arXiv:2402.06196.
- Popović (2017) Maja Popović. 2017. chrF++: words helping character n-grams. In Proceedings of the Second Conference on Machine Translation, pages 612–618, Copenhagen, Denmark. Association for Computational Linguistics.
@@
- Rei et al. (2020) Ricardo Rei, Craig Stewart, Ana C Farinha, and Alon Lavie. 2020. COMET: A neural framework for MT evaluation. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 2685–2702, Online. Association for Computational Linguistics.
- RTL Lëtzebuerg (2025) RTL Lëtzebuerg. 2025. Rtl.lu - offiziell säit vum radio télé lëtzebuerg. Accessed: 2025-10-11.
- Tanzer et al. (2023) Garrett Tanzer, Mirac Suzgun, Eline Visser, Dan Jurafsky, and Luke Melas-Kyriazi. 2023. A benchmark for learning to translate a new language from one grammar book. ArXiv, abs/2309.16575.
- Team et al. (2022a) NLLB Team, Marta R. Costa-jussà, James Cross, Onur Çelebi, Maha Elbayad, Kenneth Heafield, Kevin Heffernan, Elahe Kalbassi, Janice Lam, Daniel Licht, Jean Maillard, Anna Sun, Skyler Wang, Guillaume Wenzek, Al Youngblood, Bapi Akula, Loic Barrault, Gabriel Mejia Gonzalez, Prangthip Hansanti, John Hoffman, Semarley Jarrett, Kaushik Ram Sadagopan, Dirk Rowe, Shannon Spruit, Chau Tran, Pierre Andrews, Necip Fazil Ayan, Shruti Bhosale, Sergey Edunov, Angela Fan, Cynthia Gao, Vedanuj Goswami, Francisco Guzmán, Philipp Koehn, Alexandre Mourachko, Christophe Ropers, Safiyyah Saleem, Holger Schwenk, and Jeff Wang. 2022a. No language left behind: Scaling human-centered machine translation.
- Team et al. (2022b) NLLB Team, Marta R. Costa-jussà, James Cross, Onur Çelebi, Maha Elbayad, Kenneth Heafield, Kevin Heffernan, Elahe Kalbassi, Janice Lam, Daniel Licht, Jean Maillard, Anna Sun, Skyler Wang, Guillaume Wenzek, Al Youngblood, Bapi Akula, Loic Barrault, Gabriel Mejia Gonzalez, Prangthip Hansanti, John Hoffman, Semarley Jarrett, Kaushik Ram Sadagopan, Dirk Rowe, Shannon Spruit, Chau Tran, Pierre Andrews, Necip Fazil Ayan, Shruti Bhosale, Sergey Edunov, Angela Fan, Cynthia Gao, Vedanuj Goswami, Francisco Guzmán, Philipp Koehn, Alexandre Mourachko, Christophe Ropers, Safiyyah Saleem, Holger Schwenk, and Jeff Wang. 2022b. No language left behind: Scaling human-centered machine translation.
- (23) Jörg Tiedemann, Mikko Aulamo, Daria Bakshandaeva, Michele Boggia, Stig-Arne Grönroos, Tommi Nieminen, Alessandro Raganato, Yves Scherrer, Raúl Vázquez, and Sami Virpioja. Democratizing neural machine translation with OPUS-MT. 58(2):713–755.
- Warstadt et al. (2020) Alex Warstadt, Alicia Parrish, Haokun Liu, Anhad Mohananey, Wei Peng, Sheng-Fu Wang, and Samuel R. Bowman. 2020. BLiMP: The benchmark of linguistic minimal pairs for English. Transactions of the Association for Computational Linguistics, 8:377–392.
@@
- Xiang et al. (2021) Beilei Xiang, Changbing Yang, Yu Li, Alex Warstadt, and Katharina Kann. 2021. CLiMP: A benchmark for Chinese language model evaluation. In Proceedings of the 16th Conference of the European Chapter of the Association for Computational Linguistics: Main Volume, pages 2784–2790, Online. Association for Computational Linguistics.
@@
- Zhang et al. (2025) Chen Zhang, Mingxu Tao, Zhiyuan Liao, and Yansong Feng. 2025. MiLiC-eval: Benchmarking multilingual LLMs for China’s minority languages. In Findings of the Association for Computational Linguistics: ACL 2025, pages 11086–11102, Vienna, Austria. Association for Computational Linguistics.
@@
- Zhong et al. (2024) Tianyang Zhong, Zhenyuan Yang, Zhengliang Liu, Ruidong Zhang, Yiheng Liu, Haiyang Sun, Yi Pan, Yiwei Li, Yifan Zhou, Hanqi Jiang, et al. 2024. Opportunities and challenges of large language models for low-resource languages in humanities research. arXiv preprint arXiv:2412.04497.

### v2

@@ Abstract
Grammar refers to the system of rules that governs the structural organization and the semantic relations among linguistic units such as sentences, phrases, and words within a given language. In natural language processing, there remains a notable scarcity of grammar-focused evaluation protocols, a gap that is even more pronounced for low-resource languages. Moreover, the extent to which large language models genuinely comprehend grammatical structure especially the mapping between syntactic structures and meanings remains under debate. To investigate this issue, we propose a Grammar-Book–Guided evaluation pipeline intended to provide a systematic and generalizable framework for grammar evaluation consisting of four key stages, and in this work we take Luxembourgish as a case study. The results show a weak positive correlation between translation performance and grammatical understanding, indicating that strong translations do not necessarily imply deep grammatical competence. Larger models perform well overall due to their semantic strength but remain weak in morphology and syntax, struggling particularly with Minimal Pair tasks, while strong reasoning ability offers a promising way to enhance their grammatical understanding. The code and detailed prompts are open-sourced at : https://github.com/DobricLilujun/GBGP.git.
@@ 1. Introduction
$\spadesuit$$\spadesuit$footnotetext: These authors contributed equally to this work.In recent years, Large Language Models (LLMs) have transformed the field of Natural Language Processing (NLP), achieving remarkable progress across a variety of tasks such as text generation and translation Minaee et al. (2024). Their ability to produce coherent, and contextually relevant text has raised important discussions about whether such models merely mimic language patterns or demonstrate a deeper understanding of linguistic structures. Furthermore, while much of the early focus has been on their fluency and performance in high-resource languages such as English, French, or German, less attention has been paid to their grammatical competence in low-resource languages Li et al. (2024a).
Grammar Challenges Machine translation, in particular, is an area where LLMs are widely applied and evaluated. Success is often measured in terms of fluency and semantic adequacy compared to the reference (golden) translation; however, these metrics do not necessarily capture whether the model adheres to the explicit grammatical rules of the target languageCaglayan et al. (2020). For widely used languages with abundant training data, grammar is often implicitly learned through statistical exposure. However, the challenge becomes much more pronounced for low-resource languages, which have limited digital representation and many corner cases that are especially difficult to capture during trainingLi et al. (2024b); Zhong et al. (2024); Cahyawijaya et al. (2024). In such cases, an LLM’s ability to generate correct translations may not imply genuine grammatical understanding. In this sense, Luxembourgish presents a compelling case study in this context. As a low-resource language with fewer available corpora compared to its neighboring languages (German and French), Luxembourgish can pose difficult challenges on both translation and language understandingLothritz et al. (2022).
@@ 2. Related Work
BERT Moment in Translation LLM remain controversial regarding whether they truly understand semantics or merely mimic human speech; a potential attitude that makes one wonder whether an LLM does not amount to being just a “stochastic parrot” Bender et al. (2021). Despite these debates, their performance in language translation has become unparalleled, vastly exceeding many previous non-transformer and even encoder–decoder models Tiedemann et al. (2024); Team et al. (2022). This development constitutes a genuine “BERT moment” for translation research. Previous non-conversational models showed limited logical reasoning, and their multilingual ability was mostly attributed to abstract mechanisms like semantic mapping. In contrast, LLMs mark a major change: by asking targeted questions and running controlled comparisons, one can directly test how they understand and apply grammatical rules, giving clearer, verifiable evidence of their cross lingual abilities.
@@ 3.1. Mindsets - Why Grammar Book?
Translation enhanced by Grammar Our work is inspired by "Machine Translation from One Book" which demonstrates that it is possible to translate a language with virtually no web data (Kalamang) using a grammar book as the primary resource Tanzer et al. (2023); Kondo et al. (2024). However, Aycock et al. (2024) shows that the translation performance in this setting is primarily attributable to the parallel example sentences embedded within the grammar book, rather than the grammatical descriptions themselves. This raises the question: instead of using the grammar book for translation, can we leverage a complete grammar book to simply evaluate the grammatical understanding capabilities of LLMs? Furthermore, open-source, grammar-focused datasets remain scarce and often exhibit incomplete coverage of grammatical phenomena, with many evaluations restricted to a single minimal-pair testing paradigm whose richness is limited.
@@ 3.2. Pipeline Structure
Material Inspector: component which identifies and typifies grammatical phenomena and error patterns from the grammar document;
@@ 3.3. Models and Datasets
The Luxembourgish-to-English pairs in MIXED derived from grammar books were primarily generated by large language models (LLMs) during the extraction process, which simultaneously produced the corresponding English translations. In contrast, FLORES200 contains 1,012 relatively short sentence pairs that were manually verified. The LONG300 subset includes long Luxembourgish sentences sampled from the web, translated into English using LLMs, and subsequently verified by human reviewers.
Dataset Avg. Char Len ( ± ) # N Topics Ref. G2040 171.63 ± 21.42 2040 Grammars 2023 LONG300 429.47 ± 49.18 300 News 2025 FLORES200 125.61 ± 41.18 1012 Diverse 2022
@@ 4.2. Statistical Analysis
We conducted Spearman and Pearson correlation analyses on the performance of the four grammatical comprehension test tasks as well as on translation performance, and the results are illustrated in Figure 2. It can be observed that Tasks 1, 2, and 3 exhibit very strong correlations, primarily reflecting the ability to identify grammatical structures and their corresponding sentences. However, these three tasks show relatively low correlations with Task 4, which involves identifying sentences with correct grammatical usage. As a result, there is a weak positive correlation between performance on grammatical comprehension tasks and translation performance; however, it is insufficient to conclude that strong translation implies grammatical understanding.
In addition, we included the relatively smaller Unbabel/wmt22-cometkiwi-da (Small size) model as a supplementary evaluation to the Unbabel/wmt23-cometkiwi-da-xl (Big size) model. We found that the larger model shows higher correlations with spBLEU and ChrF++, while the smaller model’s evaluation scores are less stable. This finding led us to adopt Unbabel/wmt22-cometkiwi-da as our primary evaluation metric.
@@ 4.3. Translation Vs Grammar Understanding?
Partial mastery of grammar We take the Llama family as a case study and observe that scaling from 1B - 3B to 70B parameters improves translation performance from spBLEU to , which yields partial gains on Task 2 and Task 3 (approximately 30%), yet leaves Task 1 essentially unchanged and Task 4 still near chance-level performance. Comparable patterns emerge for the Qwen and Gemma families. Only GPT-5 achieves consistently superior results across all tasks and surpasses chance performance on Task 4. In our evaluation of Task 4, we prompted LLMs to directly generate the grammatically correct Luxembourgish sentence; however, given that grammar constitutes a strongly logical rule system, we posit that incorporating Chain-of-Thought (CoT) reasoning may enhance model performance on this task, whereas direct inference to the final option could lead the model to select responses intuitively rather than through rigorous logical deduction. Moreover, interestingly, reasoning model “Qwen3-Next-80B-A3B-Thinking” demonstrate higher performance on most grammar understanding tasks, even outperforming GPT-5, highlighting the crucial role of logical reasoning in grammatical comprehension. Moreover, increasing model size within the same family confers modest benefits to grammatical understanding - likely via enhanced reasoning - yet these gains are markedly smaller than those observed for translation.
@@ Acknowledgements
- PyMuPDF: python library for pdf and document processing. Note: https://github.com/pymupdf/PyMuPDFVersion 1.26.5, AGPL-3.0 License, accessed 2025-10-14 Cited by: §3.2.
@@
- Can llms really learn to translate a low-resource language from one grammar book?. ArXiv abs/2409.19151. External Links: Link Cited by: §3.1.
- On the dangers of stochastic parrots: can language models be too big?. In Proceedings of the 2021 ACM Conference on Fairness, Accountability, and Transparency, FAccT ’21, New York, NY, USA, pp. 610–623. External Links: ISBN 9781450383097, Link, Document Cited by: §2.
@@
- Curious case of language generation evaluation metrics: a cautionary tale. arXiv preprint arXiv:2010.13588. Cited by: §1.
@@
- Llms are few-shot in-context low-resource language learners. arXiv preprint arXiv:2403.16512. Cited by: §1.
- PaddleOCR 3.0 technical report. External Links: 2507.05595, Link Cited by: §3.2.
- How syntax specialization emerges in language models. External Links: 2505.19548, Link Cited by: §2.
@@
- Phonologie der n-tilgung im moselfränkischen (’eifeler regel’). ein beitrag zur dialektologischen prosodieforschung. pp. 29–68. External Links: ISBN 3825351564 Cited by: §4.1.
- Luxembourgish. Oxford University Press. External Links: Document, Link Cited by: §1, Table 1.
@@
- How to make the most of LLMs’ grammatical knowledge for acceptability judgments. In Proceedings of the 2025 Conference of the Nations of the Americas Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), L. Chiruzzo, A. Ritter, and L. Wang (Eds.), Albuquerque, New Mexico, pp. 7416–7432. External Links: Link, Document, ISBN 979-8-89176-189-6 Cited by: §2.
- Enhancing translation accuracy of large language models through continual pre-training on parallel data. In Proceedings of the 21st International Conference on Spoken Language Translation (IWSLT 2024), E. Salesky, M. Federico, and M. Carpuat (Eds.), Bangkok, Thailand (in-person and online), pp. 203–220. External Links: Link, Document Cited by: §3.1.
@@
- SentencePiece: a simple and language independent subword tokenizer and detokenizer for neural text processing. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, E. Blanco and W. Lu (Eds.), Brussels, Belgium, pp. 66–71. External Links: Link, Document Cited by: §3.4.
@@
- Quantifying multilingual performance of large language models across languages. CoRR. Cited by: §1.
@@
- Language ranker: a metric for quantifying llm performance across high and low-resource languages. arxiv 2024. arXiv preprint arXiv:2404.11553. Cited by: §1.
@@
- Evaluating llms’ grammatical error correction performance in learner chinese. PLOS ONE 19 (10), pp. 1–18. External Links: Document, Link Cited by: §2.
- LuxemBERT: simple and practical data augmentation in language model pre-training for luxembourgish. In 13th Language Resources and Evaluation Conference (LREC 2022), Cited by: §1.
- Large language models: a survey. arXiv preprint arXiv:2402.06196. Cited by: §1.
- ChrF++: words helping character n-grams. In Proceedings of the Second Conference on Machine Translation, O. Bojar, C. Buck, R. Chatterjee, C. Federmann, Y. Graham, B. Haddow, M. Huck, A. J. Yepes, P. Koehn, and J. Kreutzer (Eds.), Copenhagen, Denmark, pp. 612–618. External Links: Link, Document Cited by: §3.4.
@@
- COMET: a neural framework for MT evaluation. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), B. Webber, T. Cohn, Y. He, and Y. Liu (Eds.), Online, pp. 2685–2702. External Links: Link, Document Cited by: §3.4.
- RTL.lu - offiziell säit vum radio télé lëtzebuerg. RTL Group S.A. (lb). Note: Accessed: 2025-10-11 External Links: Link Cited by: Table 1.
- A benchmark for learning to translate a new language from one grammar book. ArXiv abs/2309.16575. External Links: Link Cited by: §3.1.
- No language left behind: scaling human-centered machine translation. External Links: 2207.04672, Link Cited by: §2, Table 1.
- Democratizing neural machine translation with OPUS-MT. 58 (2), pp. 713–755. External Links: ISSN 1574-0218, Link, Document Cited by: §2.
- BLiMP: the benchmark of linguistic minimal pairs for English. Transactions of the Association for Computational Linguistics 8, pp. 377–392. External Links: Link, Document Cited by: §2.
@@
- CLiMP: a benchmark for Chinese language model evaluation. In Proceedings of the 16th Conference of the European Chapter of the Association for Computational Linguistics: Main Volume, P. Merlo, J. Tiedemann, and R. Tsarfaty (Eds.), Online, pp. 2784–2790. External Links: Link, Document Cited by: §2.
@@
- MiLiC-eval: benchmarking multilingual LLMs for China’s minority languages. In Findings of the Association for Computational Linguistics: ACL 2025, W. Che, J. Nabende, E. Shutova, and M. T. Pilehvar (Eds.), Vienna, Austria, pp. 11086–11102. External Links: Link, Document, ISBN 979-8-89176-256-5 Cited by: §2.
@@
- Opportunities and challenges of large language models for low-resource languages in humanities research. arXiv preprint arXiv:2412.04497. Cited by: §1.
- PyMuPDF: python library for pdf and document processing. Note: https://github.com/pymupdf/PyMuPDFVersion 1.26.5, AGPL-3.0 License, accessed 2025-10-14 Cited by: §3.2.
- Can llms really learn to translate a low-resource language from one grammar book?. ArXiv abs/2409.19151. External Links: Link Cited by: §3.1.
- On the dangers of stochastic parrots: can language models be too big?. In Proceedings of the 2021 ACM Conference on Fairness, Accountability, and Transparency, FAccT ’21, New York, NY, USA, pp. 610–623. External Links: ISBN 9781450383097, Link, Document Cited by: §2.
- Curious case of language generation evaluation metrics: a cautionary tale. arXiv preprint arXiv:2010.13588. Cited by: §1.
- Llms are few-shot in-context low-resource language learners. arXiv preprint arXiv:2403.16512. Cited by: §1.
- PaddleOCR 3.0 technical report. External Links: 2507.05595, Link Cited by: §3.2.
- How syntax specialization emerges in language models. External Links: 2505.19548, Link Cited by: §2.
- Phonologie der n-tilgung im moselfränkischen (’eifeler regel’). ein beitrag zur dialektologischen prosodieforschung. pp. 29–68. External Links: ISBN 3825351564 Cited by: §4.1.
- Luxembourgish. Oxford University Press. External Links: Document, Link Cited by: §1, Table 1.
- How to make the most of LLMs’ grammatical knowledge for acceptability judgments. In Proceedings of the 2025 Conference of the Nations of the Americas Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), L. Chiruzzo, A. Ritter, and L. Wang (Eds.), Albuquerque, New Mexico, pp. 7416–7432. External Links: Link, Document, ISBN 979-8-89176-189-6 Cited by: §2.
- Enhancing translation accuracy of large language models through continual pre-training on parallel data. In Proceedings of the 21st International Conference on Spoken Language Translation (IWSLT 2024), E. Salesky, M. Federico, and M. Carpuat (Eds.), Bangkok, Thailand (in-person and online), pp. 203–220. External Links: Link, Document Cited by: §3.1.
- SentencePiece: a simple and language independent subword tokenizer and detokenizer for neural text processing. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, E. Blanco and W. Lu (Eds.), Brussels, Belgium, pp. 66–71. External Links: Link, Document Cited by: §3.4.
- Quantifying multilingual performance of large language models across languages. CoRR. Cited by: §1.
- Language ranker: a metric for quantifying llm performance across high and low-resource languages. arxiv 2024. arXiv preprint arXiv:2404.11553. Cited by: §1.
- Evaluating llms’ grammatical error correction performance in learner chinese. PLOS ONE 19 (10), pp. 1–18. External Links: Document, Link Cited by: §2.
- LuxemBERT: simple and practical data augmentation in language model pre-training for luxembourgish. In 13th Language Resources and Evaluation Conference (LREC 2022), Cited by: §1.
- Large language models: a survey. arXiv preprint arXiv:2402.06196. Cited by: §1.
- ChrF++: words helping character n-grams. In Proceedings of the Second Conference on Machine Translation, O. Bojar, C. Buck, R. Chatterjee, C. Federmann, Y. Graham, B. Haddow, M. Huck, A. J. Yepes, P. Koehn, and J. Kreutzer (Eds.), Copenhagen, Denmark, pp. 612–618. External Links: Link, Document Cited by: §3.4.
- COMET: a neural framework for MT evaluation. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), B. Webber, T. Cohn, Y. He, and Y. Liu (Eds.), Online, pp. 2685–2702. External Links: Link, Document Cited by: §3.4.
- RTL.lu - offiziell säit vum radio télé lëtzebuerg. RTL Group S.A. (lb). Note: Accessed: 2025-10-11 External Links: Link Cited by: Table 1.
- A benchmark for learning to translate a new language from one grammar book. ArXiv abs/2309.16575. External Links: Link Cited by: §3.1.
- No language left behind: scaling human-centered machine translation. External Links: 2207.04672, Link Cited by: §2, Table 1.
- Democratizing neural machine translation with OPUS-MT. 58 (2), pp. 713–755. External Links: ISSN 1574-0218, Link, Document Cited by: §2.
- BLiMP: the benchmark of linguistic minimal pairs for English. Transactions of the Association for Computational Linguistics 8, pp. 377–392. External Links: Link, Document Cited by: §2.
- CLiMP: a benchmark for Chinese language model evaluation. In Proceedings of the 16th Conference of the European Chapter of the Association for Computational Linguistics: Main Volume, P. Merlo, J. Tiedemann, and R. Tsarfaty (Eds.), Online, pp. 2784–2790. External Links: Link, Document Cited by: §2.
- MiLiC-eval: benchmarking multilingual LLMs for China’s minority languages. In Findings of the Association for Computational Linguistics: ACL 2025, W. Che, J. Nabende, E. Shutova, and M. T. Pilehvar (Eds.), Vienna, Austria, pp. 11086–11102. External Links: Link, Document, ISBN 979-8-89176-256-5 Cited by: §2.
- Opportunities and challenges of large language models for low-resource languages in humanities research. arXiv preprint arXiv:2412.04497. Cited by: §1.